### v1

# Spectrum radii of trees

Abstract
For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of positive numbers defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):1\leq i\leq s\}$, where $1\leq s<r$, $\beta:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ as long as $\beta>0$. We first show that there exists a tree $T$ such that its maximum degree $\Delta(T)$ is at most $r$ and its spectrum radius $\lambda(T)$ is equal to $\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. It follows that the set of spectrum radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$. Applying this conclusion, we prove that for any positive integers $r$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
AMS Subject classification: 05C05; 05C20; 05C50
Keywords: tree, maximum degree, characteristic polynomial; eigenvalue; spectrum radius
1 Introduction
All graphs considered in this article are simple and undirected. For a graph $G$, let $V(G)$, $E(G)$ and $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively. For any non-empty subset $S$ of $V(G)$, let $G[S]$ denote the subgraph of $G$ induced by $S$. For any $u\in V(G)$, let $N_{G}(u)$ (or simply $N(u)$) denote the set of neighbours of $u$ in $G$, and let $d_{G}(u)$ or (simply $d(u)$) be the size of $N_{G}(u)$, called the degree of $u$ in $G$. Let $L(G)$ denote the set of vertices $v\in V(G)$ with $d_{G}(v)=1$.
For any integer $k>0$, let $\llbracket k\rrbracket=\{1,2,\dots,k\}$. For a graph $G=(V,E)$, where $V=\{v_{i}:i\in\llbracket k\rrbracket\}$, its adjacency matrix $A(G)$ according to the vertex ordering $v_{1},v_{2},\ldots,v_{n}$ is defined to be the $0-1$ matrix $(a_{i,j})_{n\times n}$, where $a_{i,j}=1$ if and only if $v_{i}v_{j}\in E$. The characteristic polynomial of $G$, denoted by $P(G,\lambda)$, is defined to be the function $det(\lambda I_{n}-A(G))$, where $I_{n}$ is the identical matrix of size $n$. It is known that $P(G,\lambda)$ is independent of the vertex ordering $v_{1},v_{2},\dots,v_{n}$, and the roots of the equation $P(G,\lambda)=0$, called the eigenvalues of $G$, are real numbers (see [6]), and the largest eigenvalue of $G$, denoted by $\lambda(G)$, is called the spectral radius of $G$.
This spectral radius of graphs, which has been studied extensively (see a survey [17] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 8]). In this article, we focus on the study of spectral radii of trees. Let $\Lambda$ denote the set of $\lambda(T)$ for all non-trivial trees $T$ (i.e., trees of order at least $2$). It is known that $q\geq 1$ for each $q\in\Lambda$. The first aim in this article is to provide a characterization of the set $\Lambda$. For any positive number $\alpha$ and positive integer $r$, we define a set ${\mathscr{W}}_{r}(\alpha)$ of positive numbers as follows:
  1. (i).
    $\alpha\in{\mathscr{W}}_{r}(\alpha)$; and
  2. (ii).
    for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s<r$, $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ whenever $\beta>0$.
It can be verified that ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha>0$, ${\mathscr{W}}_{r}(1)=\{1\}$ for all $r\geq 2$, ${\mathscr{W}}_{2}(\sqrt{2})=\{\sqrt{2},(\sqrt{2})^{-1}\}$ and ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$. In this article, we apply the set ${\mathscr{W}}_{r}(\alpha)$ to determine if there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ for any given pair of positive numbers $\alpha$ and $r$, where $r$ is an integer.
Theorem 1.1.
For any positive number $\alpha$ and positive integer $r$, there exists a tree $T$ such that $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
For any tree $T$, $|V(T)|\geq 3$ if and only if $\Delta(T)\geq 2$. In 2003, Stevanović [16] proved that for any tree $T$ with $|V(T)|\geq 3$,
$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
By the definition of ${\mathscr{W}}_{r}(\alpha)$, we have ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r^{\prime}}(\alpha)$ whenever $r<r^{\prime}$. Thus, a characterization of $\Lambda$ follows from Theorem 1.1 and (1) directly.
Corollary 1.2.
$\Lambda$ is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)$.
By (1), $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$. We wonder if, for any $\alpha\in\Lambda$ and positive integer $r$, this condition is sufficient for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
Conjecture 1.3.
Let $\alpha\in\Lambda\setminus\{1\}$. If $r$ is an integer such that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, then there exists a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
Our second aim in this article is to apply Theorem 1.1 to prove Conjecture 1.3 for the case $\alpha=\sqrt{k}$, where $k\geq 2$ is an integer. Certainly, it includes the case $\alpha=k$ for any integer $k\geq 2$.
Theorem 1.4.
For any positive integers $k$ and $r$, where $r\geq 2$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
The remainder of this article is structured as follows. In Section 2, we examine the set ${\mathscr{W}}_{r}(\alpha)$ and establish a connection between a tree $T$ with $\lambda(T)=\alpha$ and the elements of ${\mathscr{W}}_{r}(\alpha)$. In Section 3, we introduce a family of non-trivial directed trees $T$ with a unique source $z$ and a vertex-weight function $w$, ensuring that $w(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$, for any given positive numbers $r$ and $\alpha$, where $r$ is an integer. In Sections 4 and 5, we present proofs of Theorems 1.1 and 1.4, respectively. Finally, in Section 6, we discuss open problems arising from our work.
2 The set ${\mathscr{W}}_{r}(\alpha)$
Let $\alpha$ and $r$ be positive, where $r$ is an integer. By definition, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ and $0<q\leq\alpha$ for each $q\in{\mathscr{W}}_{r}(\alpha)$. The following result tells that when $r\leq\alpha^{2}/4+1$, all numbers in ${\mathscr{W}}_{r}(\alpha)$ are at least $\frac{\alpha}{2}$.
Lemma 2.1.
If $12\leq r\leq\alpha^{2}/4+$ 1, then $q>\frac{\alpha}{2}$ for each $q\in{\mathscr{W}}_{r}(\alpha)$.
Proof. Let $\beta:=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Since $12\leq r\leq\alpha^{2}/4+$ 1, we have $\alpha>\beta\geq\frac{\alpha}{2}$.
We will prove that $q>\beta$ holds for each $q\in{\mathscr{W}}_{r}(\alpha)$. Suppose this conclusion does not hold. By the definition of ${\mathscr{W}}_{r}(\alpha)$ and the fact $\alpha>\beta$, there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $1\leq s<r$, such that $q_{i}>\beta$ for all $i\in\llbracket s\rrbracket$ and
$q:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}\leq\beta.$ (2)
However, it contradicts the following conclusion:
$q=\alpha-\sum_{i=1}^{s}q_{i}^{-1}>\alpha-(r-1)\beta^{-1}=\beta,$ (3)
where the equality $\alpha-(r-1)\beta^{-1}=\beta$ follows from the assumption $\beta=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Hence the lemma follows. $\Box$
For example, if $r=2$ and $\alpha=2$, then $r=\frac{\alpha^{2}}{4}+1$. By Lemma 2.1, $q>\frac{2}{2}=1$ for all $q\in{\mathscr{W}}_{2}(2)$. This conclusion coincides with the fact that ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$. By Theorem 1.1, Lemma 2.1 implies that there is no trees $T$ with $12\leq\Delta(T)\leq\lambda(T)^{2}/4+$ 1, which coincides with inequality (1) due to Stevanović [16].
Now we are going to introduce the Perron–Frobenius theorem, by which we develop a result for studying the numbers in ${\mathscr{W}}_{r}(\alpha)$.
The Perron–Frobenius theorem, proved by Oskar Perron [14] in 1907 and Georg Frobenius [9] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices. Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectrum radius $\lambda(G)$ of a connected graph $G$ is obtained. Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$. Then, $\lambda(G)=\lambda_{1}(G)$. We say $G$ is non-trivial if $|V(G)|\geq 2$.
Theorem 2.2 (Perron–Frobenius theorem).
Let $G$ be a connected non-trivial graph. Then $\lambda_{1}(G)>\max\{0,\lambda_{2}(G)\}$ and there is an eigenvector $\bar{x}$ of $A(G)$ corresponding to $\lambda_{1}(G)$ in which all components are positive. Furthermore, $\lambda_{1}(G)$ is the only eigenvalue of $G$ with this property.
Applying Theorem 2.2, the next conclusion follows directly. Let ${\mathbb{R}}^{+}$ be the set of positive real numbers.
Proposition 2.3.
Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi$ from $V(G)$ to ${\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
$\alpha\phi(u)=\sum_{w\in N(u)}\phi(w).$ (4)
Now we are going to establish a relation between a mapping $\phi:V(G)\rightarrow{\mathbb{R}}^{+}$ satisfying the condition in (4) and the numbers in ${\mathscr{W}}_{r}(\alpha)$.
Lemma 2.4.
Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$. If $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is a mapping satisfying the condition that for each $u\in V(T)$,
$\alpha\phi(u)=\sum_{w\in N_{T}(u)}\phi(w),$ (5)
then $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ for each edge $u_{1}u_{2}$ in $T$, where $r=\Delta(T)$. Therefore, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
Proof. Let $u_{1}u_{2}$ be a fixed edge in $T$ and let $T^{\prime}$ be the subtree of $T$ induced by $\{u_{1}\}\cup V(T_{2})$, where $T_{i}$ is the component of $T-u_{1}u_{2}$ (i.e., the subgraph of $T$ obtained by deleting edge $u_{1}u_{2}$) that contains vertex $u_{i}$. Clearly, $u_{1}\in L(T^{\prime})$ (i.e., a leaf in $T^{\prime}$). For convenience, let $T^{\prime}$ denote the directed tree obtained from $T^{\prime}$ such that $u_{1}$ is its unique source (i.e., a vertex with in-degree 00), as shown in Figure 1 (b).
We are now going to prove the following claim by induction. For any vertex $w$ in $T^{\prime}$, let $d_{T^{\prime}}(u_{1},w)$ denote the distance from $u_{1}$ to $w$ in $T^{\prime}$.
Claim 1: For any directed edge $(w_{1},w_{2})$ in $T^{\prime}$, $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$.
Let $d=\max\{d_{T^{\prime}}(u_{1},w):w\in V(T^{\prime})\}$. If $w_{2}$ is a sink of $T^{\prime}$ (i.e., $w_{2}\in L(T)$), then by condition (5), $\alpha\phi(w_{2})=\phi(w_{1})$, implying that $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha\in{\mathscr{W}}_{r}(\alpha)$. Thus, Claim 1 holds for any directed edge $(w_{1},w_{2})$ of $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w_{2})=d$.
[…]
    (a) $T$                                                                  (b)     $T^{\prime}$
Figure 1: Tree $T$ and directed tree $T^{\prime}$ in ${\mathscr{T}}$
Assume that Claim 1 holds for each directed edge $(w^{\prime}_{1},w^{\prime}_{2})$ in $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w^{\prime}_{2})\geq t$, where $2\leq t\leq d$. Now we take any directed edge $(w_{1},w_{2})$ in $T^{\prime}$ such that $w_{2}$ is not a sink of $T^{\prime}$ and $d_{T^{\prime}}(u_{1},w_{2})=t-1$. Let $N_{T}(w_{2})\setminus\{w_{1}\}=\{v_{i}:i\in\llbracket s\rrbracket\},$ where $s=d_{T}(w_{2})-1\leq r-1$. Clearly, $d_{T^{\prime}}(u_{1},v_{i})=d_{T^{\prime}}(u_{1},w_{2})+1=t$ for all $i\in\llbracket s\rrbracket$. By the inductive assumption, $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. By condition (5),
$\alpha\phi(w_{2})=\phi(w_{1})+\sum_{i=1}^{s}\phi(v_{i}),$ (6)
implying that
$\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\frac{\phi(v_{i})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\left(\frac{\phi(w_{2})}{\phi(v_{i})}\right)^{-1}.$ (7)
As $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$, by the definition of ${\mathscr{W}}_{r}(\alpha)$, we have $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Hence Claim 1 holds.
It follows from Claim 1 that $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for each edge $u_{1}u_{2}$ in $T$. Taking any edge $u_{1}u_{2}\in E(T)$, where $u_{1}\in L(T)$, we have $\alpha\phi(u_{1})=\phi(u_{2})$ by (5), implying that $\alpha^{-1}=\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Thus, this result follows. $\Box$
3 Weighted Directed trees
Let ${\mathscr{T}}$ be the set of directed trees $T$ of order at least $2$ which has a unique source. Clearly, for any $T\in{\mathscr{T}}$ with the unique source $z$, the in-degree of each vertex $v\in V(T)\setminus\{z\}$ in $T$ is $1$. For any undirected non-trivial tree and any vertex $z$ in this tree, there is one and only one orientation to obtain a directed tree with $z$ as its unique source.
For any real number $\alpha$ with $\alpha\geq 1$, let ${\mathscr{T}}_{\alpha}$ be the set of trees $T\in{\mathscr{T}}$ such that (8) defines a vertex-weight function $\omega$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$:
$\omega(v)=\left\{\begin{array}[]{ll}\alpha,&\mbox{if }od_{T}(v)=0;\\ \alpha-\sum\limits_{w\in V(T)\atop(v,w)\in E(T)}\omega(w)^{-1},&\mbox{otherwise},\end{array}\right.$ (8)
where $od_{T}(v)$ is the out-degree of $v$ in $T$ and $(v,w)\in E(T)$ represents the directed edge from $v$ to $w$ in $T$. For any $T\in{\mathscr{T}}_{\alpha}$, let $\omega_{\alpha,T}$ denote the mapping defined in (8), which is unique with respect to $\alpha$ and $T$. Clearly, $\omega_{\alpha,T}(v)>0$ for all $v\in V(T)$, unless $v$ is the unique source of $T$.
For example, if $T\in{\mathscr{T}}$ is the directed tree obtained from the star $K_{1,3}$ so that its unique source is of out-degree $1$ in $T$, it can be verified that $T$ belongs to both ${\mathscr{T}}_{2}$ and ${\mathscr{T}}_{\sqrt{3}}$, as shown in Figure 2 (a) and (b), but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, as shown in Figure 2 (c) and (d), where the number beside each vertex $v$ in $T$ is the value of $\omega_{\alpha,T}(v)$.
1111111122222222
[…]
$-\frac{9}{2}$$\frac{1}{6}$$\frac{3}{2}$$\frac{3}{2}$
[…]
(a) $T\in{\mathscr{T}}_{2}$                  (b) $T\in{\mathscr{T}}_{\sqrt{3}}$                  (c) $T\notin{\mathscr{T}}_{\frac{3}{2}}$                  (d) $T\notin{\mathscr{T}}_{\sqrt{2}}$
Figure 2: $T\in{\mathscr{T}}_{2}$, $T\in{\mathscr{T}}_{\sqrt{3}}$, but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, where $T$ is oriented from $K_{1,3}$
For non-negative numbers $r$, $\alpha$ and $q$, where $r$ is a positive integer and $\alpha\geq 1$, let ${\mathscr{T}}_{r,\alpha}(q)$ denote the set of $T\in{\mathscr{T}}_{\alpha}$ satisfying the following conditions:
  1. (i).
    $\omega_{\alpha,T}(z)=q$, where $z$ is the unique source of $T$, and
  2. (ii).
    for each $v\in V(T)$, $od_{T}(v)\leq r$ if $q=0$ and $v=z$, and $od_{T}(v)\leq r-1$ otherwise.
For example, the directed tree in Figure 2 belongs to both ${\mathscr{T}}_{3,2}(1)$ and ${\mathscr{T}}_{3,\sqrt{3}}(0)$. Two directed trees of ${\mathscr{T}}_{3,2}(\frac{1}{2})$ are also shown in Figure 3.
$\frac{1}{2}$1222
[…]
    (a)                                                                  (b)
Figure 3: Two directed trees in $\mathcal{J}_{3,2}(\frac{1}{2})$
Lemma 3.1.
Let $\alpha\geq 1$, $q\geq 0$ and $r$ be a positive integer. For any $T\in{\mathscr{T}}_{r,\alpha}(q)$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$, unless $q=0$ and $v$ is the unique source of $T$.
Proof. Since $T\in{\mathscr{T}}_{r,\alpha}(q)$, we have $T\in{\mathscr{T}}_{\alpha}$. In the definition of ${\mathscr{T}}_{\alpha}$, the condition at (8) for the mapping $\omega_{\alpha,T}$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$ implies that for any vertex $v$ in $T$, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ holds, unless $v$ is the source of $T$ and $q=0$. Hence the result follows. $\Box$
4 Proof of Theorem 1.1
Lemma 4.1.
Let $\alpha\geq 1$ and $r$ be a positive integer. If there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$.
Proof. By the definition of ${\mathscr{W}}_{r}(\alpha)$ and (8), it is clear that ${\mathscr{T}}_{r,\alpha}(q)\neq\emptyset$ for any $q\in{\mathscr{W}}_{r}(\alpha)$.
Assume that $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. For each $i\in\llbracket s\rrbracket$, let $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, $\omega_{i}$ be the mapping $\omega_{\alpha,T_{i}}$ and $z_{i}$ be the unique source of $T_{i}$. By definition, $\omega_{i}(z_{i})=q_{i}$ for all $i\in\llbracket s\rrbracket$.
Now let $T$ be the directed tree obtained from $T_{1},T_{2},\dots,T_{s}$ by adding a new vertex $z$ and adding a new directed edge $(z,z_{i})$ for each $i\in\llbracket s\rrbracket$. Let $\omega:V(T)\mapsto{\mathbb{R}}^{+}\cup\{0\}$ be the mapping defined by $\omega(z)=0$ and $\omega(w)=\omega_{i}(w)$ whenever $w\in V(T_{i})$. Since $\sum\limits_{1\leq i\leq s}q_{i}^{-1}=\alpha$, condition (8) holds for $\omega$ at $z$. For any $v\in V(T)$, if $v\in V(T_{i})$, as $\omega(w)=\omega_{i}(w)$ for all $w\in V(T_{i})$, condition (8) also holds $\omega$ at $v$.
Since $s\leq r$ and $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket s\rrbracket$, we have $\Delta(T)\leq r$, we have $T\in{\mathscr{T}}_{r,\alpha}(0)$. $\Box$
Applying the set ${\mathscr{T}}_{r,\alpha}(0)$, we can now obtain the following conclusion for proving Theorem 1.1.
Lemma 4.2.
Let $\alpha\geq 1$ and $r$ be a positive integer. For any $T\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\lambda(T)=\alpha$. If $\alpha=1$, then $T\in{\mathscr{T}}_{r,\alpha}(0)$ implies that $|V(T)|=2$, and the result holds. Now assume that $\alpha>1$.
Let $T\in{\mathscr{T}}_{r,\alpha}(0)$. By definition, $T$ is a directed tree with a unique source, say $z$. By Lemma 3.1, $\omega_{\alpha,T}(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$. Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$. Clearly, $1\leq s\leq r$ and $T_{i}\in{\mathscr{T}}_{\alpha}$ for each $i\in\llbracket s\rrbracket$. Let $z_{i}$ be the vertex in $T_{i}$ such that $(z,z_{i})$ is a directed edge in $T$. Then, $z_{i}$ the unique source of $T_{i}$ and $q_{i}=\omega_{\alpha,T}(z_{i})$ for each $i\in\llbracket s\rrbracket$. By (8), we have
$0=\omega_{\alpha,T}(z)=\alpha-\sum_{i=1}^{s}q_{i}^{-1}.$ (9)
In order to prove that $\lambda(T)=\alpha$, by Proposition 2.3, it suffices to show that the mapping $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ defined below satisfies condition (4):
  1. (i).
    $\phi(z)=1$; and
  2. (ii).
    for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$, where $\omega$ simply denotes $\omega_{\alpha,T}$.
We can first show that condition (4) is satisfied for the unique source and all sinks of $T$. As $z$ is the only source of $T$ and $N^{+}_{T}(z)=\{z_{i}:i\in\llbracket s\rrbracket\}$, where $N^{+}_{T}(z)$ is the set of out-neighbors of $z$ in $T$, we have
$\sum_{v\in N_{T}(z)}\phi(v)=\sum_{i=1}^{s}\phi(z_{i})=\sum_{i=1}^{s}\frac{\phi(z)}{\omega(z_{i})}=\phi(z)\sum_{i=1}^{s}q_{i}^{-1}=\alpha\phi(z),$ (10)
where the last step follows from (9).
For any sink $w$ of $T$, $\omega(w)=\alpha$ by (8). If $w^{\prime}$ is the unique in-neighbor of $w$ in $T$, then $\phi(w^{\prime})=\phi(w)\omega(w)$ by the definition of $\phi$, and
$\sum_{v\in N_{T}(w)}\phi(v)=\phi(w^{\prime})=\phi(w)\omega(w)=\alpha\phi(w).$ (11)
It remains to consider any vertex $w$ in $T$ which is neither the source nor a sink of $T$. Assume that $w^{\prime}$ ie the only in-neighbor of $w$ and $N^{+}_{T}(w)=\{w_{1},w_{2},\dots,w_{t}\}$. By the definition of $\phi$,
$\phi(w^{\prime})+\sum_{i=1}^{t}\phi(w_{i})=\omega(w)\phi(w)+\sum_{i=1}^{t}\frac{\phi(w)}{\omega(w_{i})}=\phi(w)\left(\omega(w)+\sum_{i=1}^{t}\omega(w_{i})^{-1}\right)=\alpha\phi(w),$ (12)
where the last step follows from (8). Thus, $\phi$ satisfies condition (4) for all vertices in $T$, implying that $\lambda(T)=\alpha$ by Proposition 2.3. $\Box$
Now we are going to establish the following conclusion, by which Theorem 1.1 follows directly.
Proposition 4.3.
For any positive numbers $\alpha$ and $r$, where $\alpha\geq 1$ and $r$ is an integer, the following three statements are pairwise equivalent:
  1. (i).
    there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$;
  2. (ii).
    $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$; and
  3. (iii).
    there exist $q_{1},q_{2},\dots,q_{s}\in{\mathscr{W}}_{r}(\alpha)$, where $1\leq s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$.
Proof. (iii) follows from (ii) directly. By Lemmas 4.2 and 4.1, (i) follows from (iii).
(i) $\Rightarrow$ (ii) Assume that $T$ is a tree with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$. By Proposition 2.3, there exists a mapping $\phi$ from $V(T)$ to ${\mathbb{R}}^{+}$ such that $\alpha\phi(u)=\sum\limits_{w\in N_{T}(u)}\phi(w)$ for each $u\in V(T)$. Then, by Lemma 2.4, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. Thus, (i) implies (ii). $\Box$
Clearly, Theorem 1.1 follows from Proposition 4.3 directly. We end this section by a result on the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
For any directed graph $D$, let $\Delta^{(o)}(D)$ denote the maximum out-degree of $D$. By Lemma 4.2 and its proof, the following conclusion holds.
Proposition 4.4.
Let $\alpha\geq 1$ and $r$ be an integer with $\alpha^{2}/4+1<r\leq\alpha^{2}$. Assume that there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s\leq r$, such that $\sum\limits_{i=1}^{s}q_{i}^{-1}=\alpha$. If $s=r$ or $\Delta^{(o)}(T_{i})=r-1$ for some $T_{i}\in{\mathscr{T}}_{r,\alpha}(q_{i})$, where $i\in\llbracket s\rrbracket$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
Corollary 4.5.
For any positive real number $\alpha$ and positive integer $r$ with $\alpha^{2}/4+1<r\leq\alpha^{2}$, if there exist a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket r\rrbracket\}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
5 Proof of Theorem 1.4
Let $k$ be an integer with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. Clearly, for any integer $r$, $r>\frac{1}{4}k+1$ if and only if $r\geq r_{0}$. If $k=4m+t$, where $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
In order to prove Theorem 1.4, by Corollary 4.5, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{r}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$. We start with the following simple result.
Lemma 5.1.
Let $k=4m+t\geq 2$, where $m$ and $t$ are integers with $m\geq 0$ and $0\leq t\leq 3$.
  1. (i).
    When $m=0$ and $1\leq t\leq 3$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[1,t]$; and
  2. (ii).
    when $m\geq 1$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[3m+t-1,4m+t]$.
Proof. (i). As $m=0$, we have $k=t$ and thus, $\sqrt{k}=\sqrt{t}$. By definition, $\frac{t}{\sqrt{t}}=\sqrt{t}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ and for $2\leq t\leq 3$,
$\frac{t-1}{\sqrt{t}}=\sqrt{t}-\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t}).$
Thus, it mains to show that $\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ for $t=3$, i.e., $\frac{1}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$. Note that $\frac{2}{\sqrt{3}}\in{\mathscr{W}}_{m+2}(\sqrt{3})$. Thus,
$\frac{\sqrt{3}}{2}=\sqrt{3}-\frac{1}{\frac{2}{\sqrt{3}}}\in{\mathscr{W}}_{m+2}(\sqrt{3}),$
and
$\frac{1}{\sqrt{3}}=\sqrt{3}-\frac{1}{\frac{\sqrt{3}}{2}}\in{\mathscr{W}}_{m+2}(\sqrt{3}).$
(ii) Assume that $m\geq 1$. For any integer $s$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
$\frac{s}{\sqrt{4m+t}}=\sqrt{4m+t}-\frac{4m+t-s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{k}).$ (13)
Thus, the result holds. $\Box$
We are now going to establish the conclusion that for any integer $k=4m+t$, where $m\geq 1$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ holds for all integers $s$ in the interval $[m,4m+t]$, on a case-by-case basis of $t$.
Proposition 5.2.
For any positive integer $m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 5.1.
Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i\geq 1$.
By definition, $2\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus, the claim holds for $i=1$. For any positive integer $i$, if $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then by definition,
$\frac{i+2}{i+1}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{i+1}{i}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (14)
Thus, the claim holds.
Claim 2: $\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ and $\frac{m-1}{m}{\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Claim 1, $\frac{m+1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus,
$\sqrt{m}=2\sqrt{m}-\frac{m+1}{\frac{m+1}{m}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (15)
and
$\frac{m-1}{m}{\sqrt{m}}=2\sqrt{m}-\frac{m+1}{\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (16)
Thus, Claim 2 holds.
Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $20\leq i\leq m-$ 2.
By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any integer $i$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
$\frac{m-i-1}{m-i}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{m-i}{m-i+1}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (17)
Thus, the claim holds.
Claim 4: For any integer $s$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Lemma 5.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any integer $i$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
$\frac{2m+i}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{(m-i-1){\sqrt{m}}}{m-i}}-\frac{i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (18)
Hence Claim 4 holds for $22m\leq s\leq 3m-$ 2. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
For any integer $i$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
$\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}=2\sqrt{m}-\frac{m+1}{\frac{(2m-i)}{2\sqrt{m}}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ (19)
and
$\frac{2m-i-1}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}}-\frac{i}{\sqrt{m}}-\frac{1}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (20)
Thus, $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all $i$ with $0\leq i\leq m$. Claim 4 holds.
By Claim 4, the conclusion holds. $\Box$
Proposition 5.3.
For any positive integer $m$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
Proof. We first prove the following claims.
Claim 1: For all $i\geq 1$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 5.1, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
$b:=\sqrt{4m+1}-\frac{m}{Q(i)}=\sqrt{4m+1}-\frac{m}{\frac{(2i+1)m}{i\sqrt{4m+1}}}=\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ (21)
and
$Q(i+1)=\frac{(2i+3)m}{(i+1)\sqrt{4m+1}}=\sqrt{4m+1}-\frac{m}{b}-\frac{1}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (22)
Thus, Claim 1 follows.
Claim 2: For any integer $i$ with $1\leq i\leq m$, we have $\frac{3m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
By Claim 1, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for $i\geq 1$. If $1\leq i\leq m$, then
$\frac{3m-i}{\sqrt{4m+1}}=\sqrt{4m+1}-\frac{i+1}{\frac{(i+1)\sqrt{4m+1}}{2i+1}}-\frac{m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (23)
Claim 2 follows.
Claim 3: For any integer $i$ with $0\leq i\leq m$, we have $\frac{2m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
Let $P(i)=\frac{2m-i}{\sqrt{4m+1}}$. By Claim 2, $P(0)=\frac{2m}{\sqrt{4m+1}}=\frac{3m-m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. By definition,
$\sqrt{4m+1}-\frac{m+1}{\frac{2m}{\sqrt{4m+1}}}=\frac{(m-1)\sqrt{4m+1}}{2m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ (24)
and thus
$P(1)=\frac{2m-1}{\sqrt{4m+1}}=\sqrt{4m+1}-\frac{m-1}{\frac{(m-1)\sqrt{4m+1}}{2m}}-\frac{2}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (25)
Hence the claim holds for $i=0,1$. It suffices to show that for any integer $i$ with $20\leq i\leq m-$ 2, if $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Assume that $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then
$b:=\sqrt{4m+1}-\frac{m+1}{P(i)}=\sqrt{4m+1}-\frac{m+1}{\frac{2m-i}{\sqrt{4m+1}}}=\frac{(m-i-1)\sqrt{4m+1}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (26)
By Claim 1, $\frac{(i+2)\sqrt{4m+1}}{2i+3}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Note that
$\sqrt{4m+1}-\frac{m-i-1}{b}-\frac{i+2}{\frac{(i+2)\sqrt{4m+1}}{2i+3}}$ $=$ $\sqrt{4m+1}-\frac{m-i-1}{\frac{(m-i-1)\sqrt{4m+1}}{2m-i}}-\frac{i+2}{\frac{(i+2)\sqrt{4m+1}}{2i+3}}$ (27)
$=$ $\frac{2m-i-2}{\sqrt{4m+1}},$
implying that $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ by definition. Thus, Claim 3 holds.
By Lemma 5.1 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 5.4.
For any positive integer $m$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
Proof. We first prove the following claims.
Claim 1: For any integer $i\geq 1$, both $\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$ belong to ${\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
Let $P(i)=\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $Q(i)=\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$. By Lemma 5.1, $P(1)=\frac{3m+1}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ and $Q(1)=\frac{3m+2}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 1 holds for $i=1$.
If $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$, then
$\sqrt{4m+2}-\frac{m+1}{\frac{(2i+1)m+i}{i\sqrt{4m+2}}}=\frac{\sqrt{4m+2}\,m(i+1)}{i(2m+1)+m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (28)
and
$Q(i+1)=\frac{(2(i+1)+1)m+(i+1)+1}{(i+1)\,\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m}{\frac{\sqrt{4m+2}\,m(i+1)}{i(2m+1)+m}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (29)
If $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$, then
$\sqrt{4m+2}-\frac{m}{\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}}=\frac{\sqrt{4m+2}\,(i+1)(m+1)}{i(2m+1)+m+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (30)
and
$P(i+1)=\frac{(2(i+1)+1)m+(i+1)}{(i+1)\,\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{\sqrt{4m+2}\,(i+1)(m+1)}{i(2m+1)+m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (31)
Thus, Claim 1 holds.
Claim 2: $\frac{2m-i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $i$ with $-2\leq i\leq m+1$.
Let $W(i)=\frac{2m-i}{\sqrt{4m+2}}$. By Claim 1, $W(-2)=\frac{(2m+1)m+m}{m\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. By definition,
$W(-1)=\frac{2m+1}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{2m+2}{\sqrt{4m+2}}}=\sqrt{4m+2}-\frac{m+1}{W(-2)}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (32)
$W(0)=\frac{2m}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{2m+1}{\sqrt{4m+2}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (33)
and
$W(1)=\frac{2m-1}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{1}{\frac{2m+1}{\sqrt{4m+2}}}-\frac{m}{\frac{2m}{\sqrt{4m+2}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (34)
Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $-2\leq i\leq 1$.
It suffices to show that for any $i$ with $10\leq i\leq m-$ 1, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Observe that for any $i$ with $10\leq i\leq m-$ 1,
$b:=\sqrt{4m+2}-\frac{m+1}{W(i)}=\sqrt{4m+2}-\frac{m+1}{\frac{2m-i}{\sqrt{4m+2}}}=\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (35)
and
$W(i+2)=\frac{2m-i-2}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m-i-1}{b}-\frac{i+2}{W(-1)},$ (36)
implying that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 2 holds.
Claim 3: $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $s$ with $2m+1\leq s\leq 3m$.
By Claim 2, $\frac{2m+i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for $i=0,1$. For any integer $i$ with $20\leq i\leq m-$ 2, it is shown in (35) that $\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ Thus,
$\frac{2m+2+i}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m-i-1}{\frac{(m-i-1)\cdot\sqrt{4m+2}}{2m-i}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (37)
Hence Claim 3 holds.
By Lemma 5.1 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 5.5.
For any positive integer $m$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.
Proof. We first prove the following claims.
Claim 1: For any positive integer $i$, $\frac{(i+1)\sqrt{4m+3}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ and $\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
Let $f_{1}(i)=\frac{(i+1)\sqrt{4m+3}}{2i+1}$ and $f_{2}(i)=\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}$. Obviously, $\sqrt{4m+3}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$,
$f_{2}(1)=\frac{3(m+1)}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ (38)
and
$f_{1}(1)=\frac{2\sqrt{4m+3}}{3}=\sqrt{4m+3}-\frac{m+1}{\frac{3(m+1)}{\sqrt{4m+3}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (39)
Thus, Claim 1 holds for $i=1$.
Now assume that $t$ is an integer with $t\geq 2$ such that $f_{1}(i),f_{2}(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $i$ with $11\leq i\leq t-$ 1. Then, by definition,
$\frac{2mi+i+m}{i\sqrt{4m+3}}=\sqrt{4m+3}-\frac{i+1}{\frac{(i+1)\sqrt{4m+3}}{2i+1}}-\frac{m-i}{\frac{i\sqrt{4m+3}}{2i-1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (40)
$\sqrt{4m+3}-\frac{m+1}{\frac{2mi+i+m}{i\sqrt{4m+3}}}=\frac{(i+1)m\sqrt{4m+3}}{2mi+i+m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (41)
$f_{2}(i+1)=\frac{(2i+3)(m+1)}{(i+1)\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m}{\frac{(i+1)m\sqrt{4m+3}}{2mi+i+m}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (42)
and
$f_{1}(i+1)=\frac{(i+2)\sqrt{4m+3}}{2i+3}=\sqrt{4m+3}-\frac{m+1}{f_{2}(i+1)}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (43)
Thus, Claim 1 holds.
Claim 2: For any integer $i$ with $-3\leq i\leq m$, $\frac{2m-i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
Let $g(i)=\frac{2m-i}{\sqrt{4m+3}}$. By Claim 1, $g(-3)=\frac{2m+3}{\sqrt{4m+3}}=f_{2}(m+1)$. By definition,
$\sqrt{4m+3}-\frac{m+1}{f_{1}(m)}=\sqrt{4m+3}-\frac{m+1}{\frac{(m+1)\sqrt{4m+3}}{2m+1}}=\frac{2(m+1)}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (44)
$\sqrt{4m+3}-\frac{m+1}{\frac{2(m+1)}{\sqrt{4m+3}}}=\frac{\sqrt{4m+3}}{2}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (45)
$\sqrt{4m+3}-\frac{m+1}{\frac{\sqrt{4m+3}}{2}}=\frac{2m+1}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (46)
$\sqrt{4m+3}-\frac{m+1}{\frac{2m+1}{\sqrt{4m+3}}}=\frac{m\sqrt{4m+3}}{2m+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (47)
and
$\sqrt{4m+3}-\frac{m}{\frac{m\sqrt{4m+3}}{2m+1}}-\frac{1}{\frac{\sqrt{4m+3}}{2}}=\frac{2m}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (48)
implying that $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each $i\in\{-2,-1,0\}$.
It remains to show that for any integer $i$ with $10\leq i\leq m-$ 1, $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ implies that $g(i+1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Observe that
$\sqrt{4m+3}-\frac{m+1}{g(i)}=\sqrt{4m+3}-\frac{(m+1)}{\frac{2m-i}{\sqrt{4m+3}}}=\frac{(m-i-1)\sqrt{4m+3}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (49)
By (45), $\frac{\sqrt{4m+3}}{2}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Thus,
$g(i+1)=\frac{2m-i-1}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m-i-1}{\frac{(m-i-1)\sqrt{4m+3}}{2m-i}}-\frac{i+2}{\frac{\sqrt{4m+3}}{2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (50)
Hence Claim 2 follows.
Claim 3: For any integer $s$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
By Lemma 5.1, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $33m+2\leq s\leq 4m+$ 3.
Let $h(i)=\frac{2m+i}{\sqrt{4m+3}}$. By Claim 2, $h(0),h(1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. For any integer $i$ with $0\leq i\leq m$, by Claim 1, $\frac{(m-i+1)\sqrt{4m+3}}{2(m-i)+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$, implying that
$h(i+2)=\frac{2m+2+i}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{(m+1-i)}{\frac{(m-i+1)\sqrt{4m+3}}{2\cdot(m-i)+1}}-\frac{i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (51)
Thus, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $12m\leq s\leq 3m+$ 1. Hence Claim 3 holds. The result then follows from Claims 2 and 3. $\Box$
We are now going to prove Theorem 1.4.
Proof of Theorem 1.4: The necessity follows from (1). Now we prove the sufficiency.
Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. By Lemma 5.1 and Propositions 5.25.35.4 and 5.5, for each integer $s$ with $\left\lfloor{\frac{1}{4}k}\right\rfloor\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ holds, implying that
$s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}.$ (52)
Then, by Corollary 4.5, for any integer $r$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$. $\Box$
6 Problems and Remarks
Recall that $\Lambda$ is the set of $\lambda(T)$ over all trees $T$ of order at least $2$. Conjecture 1.3 asks an interesting question on the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$ for any given $\alpha\in\Lambda\setminus\{1\}$ and integer $r$ satisfying $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
Another interesting problem is on the numbers contained in $\Lambda$. Let ${\cal P}$ be the set of real-rooted monic polynomials with integer coefficients. Salez [15] showed that every totally real algebraic integer (i.e. a root of some polynomial in ${\cal P}$) is an eigenvalue of some finite tree. Now let $\Lambda^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
Problem 6.1.
Is there a positive real number $c$ such that $\{\alpha\in\Lambda^{\prime}:\alpha\geq c\}\subseteq\Lambda$?
It can be proved by applying Proposition 4.4 that, for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{2(2p+1)+2\sqrt{4q+1}}}{2}\in\Lambda$ if $q\leq p^{2}+p$. As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ whenever $4|(p-q)$.
Proposition 6.2.
For any positive integers $p$ and $q$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\{q,\frac{p-q}{4}\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
Proof. Assume that $p-q=4d$. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
$\sqrt{q}=\frac{\sqrt{p}+\sqrt{q}}{2}-\frac{d}{\frac{\sqrt{p}+\sqrt{q}}{2}}\in{\mathscr{W}}_{r}(\alpha)$ (53)
and
$\frac{\sqrt{p}-\sqrt{q}}{2}=\frac{\sqrt{p}+\sqrt{q}}{2}-\frac{q}{\sqrt{q}}\in{\mathscr{W}}_{r}(\alpha).$ (54)
Observe that
$d\cdot\left(\frac{\sqrt{p}-\sqrt{q}}{2}\right)^{-1}=\frac{\frac{p-q}{4}}{\frac{\sqrt{p}-\sqrt{q}}{2}}=\frac{\sqrt{p}+\sqrt{q}}{2}=\alpha.$ (55)
By Proposition 4.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
We wonder if $4|(p-q)$ is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\Lambda$. For any positive integers $s$ and $t$, $\frac{s+t}{2}=\frac{\sqrt{s^{2}}+\sqrt{t^{2}}}{2}$. As $P(G,\lambda)\in{\cal P}$, each rational root of $P(G,\lambda)=0$ is an integer. It follows that $\frac{s+t}{2}\in\Lambda$ if and only if $s$ and $t$ have the same parity, i.e., $4|(s^{2}-t^{2})$.
Problem 6.3.
Is it true that for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $p-q$ is a multiple of $4$?
Let $T$ be any tree with $\alpha=\lambda(T)$ and $r=\Delta(T)$. Then, (1) yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$. By Theorem 1.1, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
Problem 6.4.
Let $\lambda\in\Lambda$ and $r$ be an integer with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
Acknowledgement
This research is supported by NSFC (No. 12101347 and 12371340), NSF of Shandong Province (No. ZR2021QA085) and the Ministry of Education, Singapore, under its Academic Research Tier 1 (RG19/22). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.