### v3

# Existence of trees with prescribed maximum degrees and spectral radii

Abstract
It is well known that the spectral radius $\rho(T)$ of a tree $T$ with at least $3$ vertices has the property that $\frac{1}{4}\rho(T)^{2}+1<\Delta(T)\leq\rho(T)^{2}$, where $\Delta(T)$ is the maximum degree of $T$. Let $\mathbb{P}$ denote the set of spectral radii of all non-trivial trees. In this article, we study the inverse problem that for any $\alpha\in\mathbb{P}$ and integer $r$ satisfying the condition $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, is there a tree $T$ such that $\Delta(T)=r$ and $\rho(T)=\alpha$?
For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote a set of non-negative real numbers defined as follows: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,1\leq i\leq s\}$, if $\beta:=\alpha-\sum\limits_{i=1}^{s}q_{i}^{-1}\geq 0$ and $s\leq r-\left\lceil\frac{\beta}{\beta+1}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$. We first show that $0\in{\mathscr{W}}_{r}(\alpha)$ if and only if there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$. It follows directly that $\mathbb{P}$ is exactly the set of positive numbers $\alpha$ such that $0\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$. Applying this conclusion, we prove that for any two positive integers $r\geq 2$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
AMS Subject classification: 05C05; 05C20; 05C50
Keywords: tree, maximum degree, characteristic polynomial; eigenvalue; spectral radius
1 Introduction
All graphs in this article are finite, simple and undirected. For a graph $G$, let $V(G)$, $E(G)$ $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively. If $|V(G)|=n$, then $G$ has exactly $n$ eigenvalues, and they are all real numbers (see [6]). The largest eigenvalue of $G$, denoted by $\rho(G)$, is called the spectral radius of $G$.
The spectral radii of graphs, which have been studied extensively (see a survey [33] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 11, 18, 19, 29, 30, 33]). The distribution of spectral radii among different graphs is a topic of ongoing research with connections to various graph properties and invariants (see [7, 9, 16, 18, 19, 23, 27, 29, 30]).
In this article, we focus on the study of spectral radii of trees. There are various known results on the relation between $\rho(T)$ and other parameters of a tree $T$. Lovász and Pelikán [22] determined the maximal and minimal value of $\rho(T)$ over all trees $T$ with a fixed order, Guo and Tan [13] found an upper bound of $\rho(T)$ in terms of the order and matching number of $T$, and Simić and Tos̆ić [31] confirmed the tree that has the largest spectral radius over all trees with given order and maximum degree. Stevanović [32] provided a relation between $\rho(T)$ and $\Delta(T)$, as stated below.
Theorem 1 ([32]).
For any tree $T$ with $|V(T)|\geq 3$, $\frac{1}{4}\rho(T)^{2}+1<\Delta(T)\leq\rho(T)^{2}.$
A graph $G$ is said to be non-trivial if $|V(G)|\geq 2$. Let $\mathbb{P}$ denote the set of $\rho(T)$ over all non-trivial trees $T$. In this article, we will study the following inverse problem:
Problem 1.
Is it true that for any $\alpha\in\mathbb{P}$ and any integer $r$ satisfying the condition $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, there exists a tree $T$ with $\rho(T)=\alpha$ and $\Delta(T)=r$?
The above inverse problem connects to characterizing possible spectra of graphs and is similar in spirit to the inverse eigenvalue problem for general matrices. Studying inverse problems in graph theory is important because it offers a reverse perspective on fundamental questions, connects structure to invariants, underpins classical existence theorems, inspires constructive algorithms, etc. In fact, many famous problems in graph theory are inverse in nature. For example, the reconstruction conjecture in graph theory (see [34]) says that any connected graph with at least three vertices is determined uniquely by its subgraphs obtained by removing one vertex, the realization problem of degree sequences studies sequences of positive integers which are degree sequences of graphs (see Erdős and Gallai [10], Havel [15] and Hakimi [14]), the inverse spectral problem studies possible sets of real numbers which are spectra of graphs (see [17]), and the inverse Wiener problem studies whether a given integer can be the Wiener index of some graph (see [20, 35, 36]).
Let $\mathbb{N}$$^{+}$ and $\mathbb{R}$$^{+}$ denote the set of positive integers and the set of positive real numbers, respectively. For any $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of non-negative real numbers defined below:
  1. (i).
    $\alpha\in{\mathscr{W}}_{r}(\alpha)$; and
  2. (ii).
    for any $s\in{\mathbb{N}}^{+}$ and multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket s\rrbracket\big{\}}$, if $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}\geq 0$ and $1\leq s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$.
It can be verified that ${\mathscr{W}}_{1}(1)=\{0,1\}$, ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha\in{\mathbb{R}}^{+}$ with $\alpha\neq 1$, ${\mathscr{W}}_{2}(\sqrt{2})=\left\{0,\sqrt{2},(\sqrt{2})^{-1}\right\}$ and ${\mathscr{W}}_{2}(2)=\big{\{}\frac{i+1}{i}:i\geq 1\big{\}}$. Clearly, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ holds for any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$.
Our first aim in this article is to establish the following result, providing a connection between the existence of a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$ and whether $0\in{\mathscr{W}}_{r}(\alpha)$.
Theorem 2.
For any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$, there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
A characterization of the set $\mathbb{P}$ follows directly from Theorems 1 and 2.
Corollary 1.
$\mathbb{P}=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)\big{\}}$.
If $r$ is removed from the definition of the set ${\mathscr{W}}_{r}(\alpha)$, we can define the set ${\mathscr{W}}(\alpha)$ as follows:
  1. (i).
    $\alpha\in{\mathscr{W}}(\alpha)$; and
  2. (ii).
    for any multi-set $\big{\{}q_{i}\in{\mathscr{W}}(\alpha):q_{i}>0,i\in\llbracket s\rrbracket\big{\}}$, if $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}\geq 0$, then $\beta\in{\mathscr{W}}(\alpha)$.
Clearly, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}(\alpha)$ for any $r\in{\mathbb{N}}^{+}$, and the following conclusion holds.
Corollary 2.
$\mathbb{P}=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}(\alpha)\big{\}}$.
Our second aim in this article is to apply Theorem 2 to prove that Problem 1 has a positive answer for the case $\alpha=\sqrt{k}$ for any $k\in{\mathbb{N}}^{+}\setminus\{1\}$. Certainly, it includes the case $\alpha=k\in{\mathbb{N}}^{+}\setminus\{1\}$.
Theorem 3.
For any $k,r\in{\mathbb{N}}^{+}\setminus\{1\}$, there exists a tree $T$ with $\rho(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
The remainder of this article is structured as follows. In Sections 2 and 3, we prove Theorem 2's necessity and sufficiency, respectively. In Section 4, we complete the proof of Theorem 3. Finally, in Section 5, we provide some open problems arising from our work.
2 The necessity of Theorem 2
The Perron–Frobenius theorem, proved by Oskar Perron [26] in 1907 and Georg Frobenius [12] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices. Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectral radius $\rho(G)$ of a connected graph $G$ is obtained. Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$. Then, $\rho(G)=\lambda_{1}(G)$.
Theorem 4 (Perron–Frobenius theorem).
Let $G$ be a connected non-trivial graph. Then $\lambda_{1}(G)>\max\{0,\lambda_{2}(G)\}$ and there is an eigenvector $\bar{x}$ of $A(G)$ corresponding to $\lambda_{1}(G)$ in which all components are positive. Furthermore, $\lambda_{1}(G)$ is the only eigenvalue of $G$ with this property.
Applying Theorem 4, the next conclusion follows directly.
Proposition 1.
Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\rho(G)=\alpha$ if and only if there is a mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
$\alpha\phi(u)=\sum_{w\in N(u)}\phi(w).$ (1)
Note that for any mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$, the condition of (1) holds if and only if $\phi$ is an eigenvector of $A(G)$ with respect to $\rho(G)$. Now we are going to prove the following lemma by which the necessity of Theorem 2 follows directly.
Lemma 1.
Let $T$ be a non-trivial tree and $\alpha=\rho(T)$, and let $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ be an eigenvector of $A(T)$ with respect to $\alpha$. Then both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$, where $r=\Delta(T)$.
Proof. Since $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is an eigenvector of $A(T)$ with respect to $\alpha=\rho(T)$, (1) holds for each $u\in V(T)$. Let's first prove the following claim.
Claim 1: For an arbitrary vertex $z$ in $T$, if $T^{\prime}$ is the orientation of $T$ with $z$ as its source, as shown in Figure 1, then $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for every directed edge $(w_{1},w_{2})$ in $T^{\prime}$.
[…]
Figure 1: A directed tree $T^{\prime}$ with $z$ as its unique source
We will prove Claim 1 by induction. For any vertex $w$ in $T^{\prime}$, let $d_{T^{\prime}}(z,w)$ denote the distance from $z$ to $w$ in $T^{\prime}$, and let $d=\max\{d_{T^{\prime}}(z,w):w\in V(T^{\prime})\}$.
If $w_{2}$ is a sink of $T^{\prime}$, then $w_{2}$ is a leaf of $T$, and the condition of (1) yields that $\alpha\phi(w_{2})=\phi(w_{1})$, implying that $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha\in{\mathscr{W}}_{r}(\alpha)$. Thus, Claim 1 holds whenever $w_{2}$ is a sink of $T^{\prime}$. Clearly, if $d_{T^{\prime}}(u_{1},w_{2})=d$, then $w_{2}$ is a sink of $T^{\prime}$ and thus Claim 1 also holds.
Assume that Claim 1 holds for each directed edge $(w^{\prime}_{1},w^{\prime}_{2})$ in $T^{\prime}$ with $d_{T^{\prime}}(z,w^{\prime}_{2})\geq t$, where $2\leq t\leq d$. Now we take any directed edge $(w_{1},w_{2})$ in $T^{\prime}$ such that $w_{2}$ is not a sink of $T^{\prime}$ and $d_{T^{\prime}}(z,w_{2})=t-1$. Assume that $N_{T}(w_{2})\setminus\{w_{1}\}=\{v_{i}:i\in\llbracket s\rrbracket\},$ where $s=d_{T}(w_{2})-1\leq r-1$. Clearly, for each $i\in\llbracket s\rrbracket$, $(w_{2},v_{i})$ is a directed edge in $T^{\prime}$ and $d_{T^{\prime}}(z,v_{i})=d_{T^{\prime}}(z,w_{2})+1=t$. By the inductive assumption, $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. By the condition of (1),
$\alpha\phi(w_{2})=\phi(w_{1})+\sum_{i=1}^{s}\phi(v_{i}),$ (2)
implying that
$\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\frac{\phi(v_{i})}{\phi(w_{2})}=\alpha-\sum_{i=1}^{s}\left(\frac{\phi(w_{2})}{\phi(v_{i})}\right)^{-1}.$ (3)
As $s<r$, by the definition of ${\mathscr{W}}_{r}(\alpha)$, we have $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Hence Claim 1 holds.
Since vertex $z$ in Claim 1 is an arbitrary vertex in $T$, Claim 1 implies that both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$. $\Box$
Proof of the necessity of Theorem 2. We follow the conditions in Lemma 1. Take any edge $u_{1}u_{2}\in E(T)$, where $u_{1}$ is a leaf of $T$. We have $\alpha\phi(u_{1})=\phi(u_{2})$ by (1), implying that $\alpha^{-1}=\frac{\phi(u_{2})}{\phi(u_{1})}\in{\mathscr{W}}_{r}(\alpha)$ by Lemma 1. Thus, $0=\alpha-\frac{1}{\alpha^{-1}}\in{\mathscr{W}}_{r}(\alpha)$ by definition. $\Box$
3 The sufficiency of Theorem 2
In this section, we always assume that $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$. Let ${\mathscr{T}}_{r,\alpha}$ denote the set of ordered pairs $(T,\omega)$, where $T$ is a directed tree with a unique source $z$ and $\omega$ is a mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ such that for each $v\in V(T)$,
  1. (i).
    $od_{T}(v)\leq r$ if $v=z$ and $\omega(z)=0$, and $od_{T}(v)\leq r-1$ otherwise; and
  2. (ii).
    $\omega(v)>0$ whenever $v\in V(T)\setminus\{z\}$, and
    $\omega(v)=\left\{\begin{array}[]{ll}\alpha,&\mbox{if }N^{+}_{T}(v)=\emptyset;\\ \alpha-\sum\limits_{u\in N^{+}_{T}(v)}\omega(u)^{-1},&\mbox{otherwise},\end{array}\right$. (4)
where $N^{+}_{T}(v)$ is the set of out-neighbors of $v$ in $T$ and $od_{T}(v)=|N^{+}_{T}(v)|$.
For any $\beta\in{\mathbb{R}}^{+}$, let ${\mathscr{T}}_{r,\alpha}(\beta)$ denote the set of ordered pairs $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}$ such that $\omega(z)=\beta$, where $z$ is the unique source of $T$. By the definitions of ${\mathscr{W}}_{r}(\alpha)$ and ${\mathscr{T}}_{r,\alpha}(\beta)$, the following conclusion holds.
Lemma 2.
For any $\beta\in{\mathscr{W}}_{r}(\alpha)$, ${\mathscr{T}}_{r,\alpha}(\beta)\neq\emptyset$.
Proof. The conclusion is trivial if $\beta=\alpha$. Now assume that $\beta=\alpha-\sum\limits_{i=1}^{s}\beta_{i}^{-1}$ and $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$ for each $i\in\llbracket s\rrbracket$, where $s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$. Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{s}$ by adding a new vertex $z$ and a new directed edge $(z,z_{i})$ for each $i\in\llbracket s\rrbracket$, where $z_{i}$ is the unique source of $T_{i}$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ such that $\omega(v)=\omega_{i}(v)$ holds for each $i\in\llbracket s\rrbracket$ and $v\in V(T_{i})$, and
$\omega(z)=\alpha-\sum\limits_{i=1}^{s}\omega(z_{i})^{-1}=\alpha-\sum\limits_{i=1}^{s}\beta_{i}^{-1}=\beta.$ (5)
Obviously, $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(\beta)$. Hence the lemma holds. $\Box$
For example, two members $(T_{i},\omega_{i})$ for $i=1,2$ of $\mathcal{J}_{3,2}(\frac{1}{2})$ are shown in Figure 2, where the value of $\omega_{i}(v)$ for each vertex $v$ in $T_{i}$ is shown beside vertex $v$.
$\frac{1}{2}$1222
$\frac{1}{2}$$\frac{2}{3}$$\frac{3}{2}$$\frac{3}{2}22$
    (a) $(T_{1},\omega_{1})$                 (b) $(T_{2},\omega_{2})$
Figure 2: $(T_{i},\omega_{i})\in\mathcal{J}_{3,2}(\frac{1}{2})$ for $i=1,2$
By Lemma 2, if $0\in{\mathscr{W}}_{r}(\alpha)$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$, and there exists $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}(0)$.
Lemma 3.
For any $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\rho(T)=\alpha$.
Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\rho(T)=\alpha$.
By definition again, $T$ is a directed tree with a unique source $z$ and $\omega(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$. Let $T_{1},T_{2},\dots,T_{s}$ be the components of TzT-z, where $s=od_{T}(z)$. Clearly, $1\leq s\leq r$. For each $i\in\llbracket s\rrbracket$, $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$, where $\omega_{i}$ is the restriction of $\omega$ to $V(T_{i})$, $\beta_{i}=\omega(z_{i})$ and $z_{i}$ is the vertex in $T_{i}$ with $z_{i}\in N^{+}_{T}(z)$. Then $z_{i}$ the unique source of $T_{i}$ for each $i\in\llbracket s\rrbracket$ and
$0=\omega(z)=\alpha-\sum_{i=1}^{s}\omega(z_{i})^{-1}=\alpha-\sum_{i=1}^{s}\beta_{i}^{-1}.$ (6)
Now let $\phi$ be the mapping from $V(T)$ to $\mathbb{R}$$^{+}$ defined below:
  1. (i).
    $\phi(z)=1$; and
  2. (ii).
    for each directed edge $(w_{1},w_{2})\in E(T)$, $\phi(w_{2})=\frac{\phi(w_{1})}{\omega(w_{2})}$.
In order to show that $\rho(T)=\alpha$, by Proposition 1, it suffices to prove the following Claim.
Claim A: $\phi$ satisfies the condition of (1) for each vertex $u\in V(T)$.
We first prove Claim A for the case $u=z$ or $u$ is a sink of $T$ (i.e., $N^{+}_{T}(u)=\emptyset$). As $N^{+}_{T}(z)=\{z_{i}:i\in\llbracket s\rrbracket\}$, we have
$\sum_{v\in N^{+}_{T}(z)}\phi(v)=\sum_{i=1}^{s}\phi(z_{i})=\sum_{i=1}^{s}\frac{\phi(z)}{\omega(z_{i})}=\phi(z)\sum_{i=1}^{s}\beta_{i}^{-1}=\alpha\phi(z),$ (7)
where the last step follows from (6).
For the case that $u$ is a sink of $T$, $\omega(u)=\alpha$ by (4). If $u^{\prime}$ is the unique in-neighbor of $u$ in $T$, then $\phi(u^{\prime})=\phi(u)\omega(u)$ by the definition of $\phi$, and
$\sum_{v\in N_{T}(u)}\phi(v)=\phi(u^{\prime})=\phi(u)\omega(u)=\alpha\phi(u).$ (8)
Finally, we assume that $u$ is vertex in $T$ which is neither the source nor a sink of $T$. Assume that $u^{\prime}$ ie the only in-neighbor of $u$ in $T$ and $N^{+}_{T}(w)=\{u_{1},u_{2},\dots,u_{t}\}$. By the definition of $\phi$,
$\phi(u^{\prime})+\sum_{i=1}^{t}\phi(u_{i})=\omega(u)\phi(u)+\sum_{i=1}^{t}\frac{\phi(u)}{\omega(u_{i})}=\phi(u)\left(\omega(u)+\sum_{i=1}^{t}\omega(u_{i})^{-1}\right)=\alpha\phi(u),$ (9)
where the last step follows from (4). Thus, $\phi$ satisfies condition (1) for each $u\in V(T)$, implying that $\rho(T)=\alpha$ by Proposition 1. $\Box$
By Lemmas 2 and 3, if $0\in{\mathscr{W}}_{r}(\alpha)$, then there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$, and thus the sufficiency of Theorem 2 holds.
Corollary 3.
If there exist a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r\rrbracket\big{\}}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\alpha$.
Proof. Assume that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, where $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ and $q_{i}>0$ for each $i\in\llbracket r\rrbracket$. By Lemma 2, there exists $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket r\rrbracket$. Assume that $z_{i}$ is the only source of $T_{i}$ for each $i\in\llbracket r\rrbracket$. Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{r}$ by adding a new vertex $z$ and directed edges $(z,z_{i})$ for each $i\in\llbracket r\rrbracket$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ defined as follows: $\omega(z)=0$ and $\omega(v)=\omega_{i}(v)$ whenever $v\in V(T_{i})$, where $i\in\llbracket r\rrbracket$. By the definition of $T$ and $\omega$, it can be verified easily that $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$. Clearly, $r\geq\Delta(T)\geq od_{T}(z)=r$, implying that $\Delta(T)=r$. By Lemma 3, $\rho(T)=\alpha$. $\Box$
Let ${\mathscr{W}}_{r,\alpha}(0)=\big{\{}\omega(v):v\in V(T),(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)\big{\}}$.
Corollary 4.
Assume that $0\in{\mathscr{W}}_{r}(\alpha)$. If ${\mathscr{W}}_{r,\alpha}(0)$ contains $\beta$ such that $\beta=\alpha-\sum\limits_{i=1}^{r-1}q_{i}^{-1}>0$ holds for a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r-1\rrbracket\big{\}}$, then there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\alpha$.
4 Proof of Theorem 3
Let $k\in{\mathbb{N}}^{+}$ with $k\geq 2$ and $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. Clearly, for any $r\in{\mathbb{N}}^{+}$, $r>\frac{1}{4}k+1$ if and only if $r\geq r_{0}$. If $k=4m+t$, where $m\in\{0\}\cup{\mathbb{N}}^{+}$ and $0\leq t\leq 3$, then $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2=m+2$.
In order to prove Theorem 3, by Corollary 3, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{s}(\sqrt{k})$) for all integers $s$ in the interval $r_{0},k]$, since this implies that $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}$. We start with the following simple result.
Lemma 4.
Let $k=4m+t\geq 2$, where $m$ and $t$ are integers with $m\geq 0$ and $0\leq t\leq 3$.
  1. (i).
    When $m=0$ and $2\leq t\leq 3$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[2,t]$; and
  2. (ii).
    when $m\geq 1$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval [3m+t1,4m+t][3m+t-1,4m+t].
Proof. (i). As $m=0$, we have $k=t$ and thus, $\sqrt{k}=\sqrt{t}$. By definition, $\frac{t}{\sqrt{t}}=\sqrt{t}\in{\mathscr{W}}_{m+2}(\sqrt{t})$ and for $2\leq t\leq 3$,
$\frac{t-1}{\sqrt{t}}=\sqrt{t}-\frac{1}{\sqrt{t}}\in{\mathscr{W}}_{m+2}(\sqrt{t}).$
(ii) Assume that $m\geq 1$. For any $s\in{\mathbb{N}}^{+}$ with $3m+t-1\leq s\leq 4m+t$, we have $0\leq 4m+t-s\leq m+1$ and
$\frac{s}{\sqrt{4m+t}}=\sqrt{4m+t}-\frac{4m+t-s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{k}).$ (10)
Thus, the result holds. $\Box$
We are now going to establish the conclusion that for any $k=4m+t$, where $m\in{\mathbb{N}}^{+}$ and $t\in\{0,1,2,3\}$, $\frac{s}{\sqrt{4m+t}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+t})$ holds for all integers $s$ in the interval $[m,4m+t]$ by considering each possible value of $t$ separately.
Proposition 2.
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $s$ in $[m,4m]$.
Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 4.
Claim 1: $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all $i\in{\mathbb{N}}^{+}$.
By definition, $2\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus, the claim holds for $i=1$. For any $i\in{\mathbb{N}}^{+}$, if $\frac{i+1}{i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then by definition,
$\frac{i+2}{i+1}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{i+1}{i}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (11)
Thus, the claim holds.
Claim 2: $\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ and $\frac{m-1}{m}{\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Claim 1, $\frac{m+1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. Thus,
$\sqrt{m}=2\sqrt{m}-\frac{m+1}{\frac{m+1}{m}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (12)
and
$\frac{m-1}{m}{\sqrt{m}}=2\sqrt{m}-\frac{m+1}{\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (13)
Thus, Claim 2 holds.
Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $0\leq i\leq m-2$.
By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any $i\in{\mathbb{N}}^{+}$ with $1\leq i\leq m-2$, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
$\frac{m-i-1}{m-i}\sqrt{m}=2\sqrt{m}-\frac{m}{\frac{m-i}{m-i+1}\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (14)
Thus, the claim holds.
Claim 4: For $s\in{\mathbb{N}}^{+}$ with $m\leq s\leq 4m$, $\frac{s}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
By Lemma 4, Claim 4 holds for $3m-1\leq s\leq 4m$. For any $i\in{\mathbb{N}}^{+}$ with $0\leq i\leq m-2$, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
$\frac{2m+i}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{(m-i-1){\sqrt{m}}}{m-i}}-\frac{i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (15)
Hence Claim 4 holds for $2m\leq s\leq 3m-2$. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
For any $i\in{\mathbb{N}}^{+}$ with $0\leq i\leq m-1$, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
$\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}=2\sqrt{m}-\frac{m+1}{\frac{(2m-i)}{2\sqrt{m}}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ (16)
and
$\frac{2m-i-1}{2{\sqrt{m}}}=2\sqrt{m}-\frac{m-i-1}{\frac{2{\sqrt{m}}\,(m-i-1)}{2m-i}}-\frac{i}{\sqrt{m}}-\frac{1}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m}).$ (17)
Thus, $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all $i$ with $0\leq i\leq m$. Claim 4 holds.
By Claim 4, the conclusion holds. $\Box$
Proposition 3.
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for all integers $s$ in $[m,4m+1]$.
Proof. We first prove the following claims.
Claim 1: For any $i\in{\mathbb{N}}^{+}$, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ and $\frac{(2i+1)m}{i\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$.
Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 4, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
$b:=\sqrt{4m+1}-\frac{m}{Q(i)}=\sqrt{4m+1}-\frac{m}{\frac{(2i+1)m}{i\sqrt{4m+1}}}=\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ (18)
and
$Q(i+1)=\frac{(2i+3)m}{(i+1)\sqrt{4m+1}}=\sqrt{4m+1}-\frac{m}{b}-\frac{1}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (19)
Thus, Claim 1 follows.
Claim 2: For any $i\in\llbracket m\rrbracket$, we have $\frac{3m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
By Claim 1, $\frac{(i+1)\sqrt{4m+1}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ for $i\geq 1$. If $1\leq i\leq m$, then
$\frac{3m-i}{\sqrt{4m+1}}=\sqrt{4m+1}-\frac{i+1}{\frac{(i+1)\sqrt{4m+1}}{2i+1}}-\frac{m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (20)
Claim 2 follows.
Claim 3: For any $i\in\{0\}\cup\llbracket m\rrbracket$, we have $\frac{2m-i}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$
Let $P(i)=\frac{2m-i}{\sqrt{4m+1}}$. By Claim 2, $P(0)=\frac{2m}{\sqrt{4m+1}}=\frac{3m-m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. By definition,
$\sqrt{4m+1}-\frac{m+1}{\frac{2m}{\sqrt{4m+1}}}=\frac{(m-1)\sqrt{4m+1}}{2m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ (21)
and thus
$P(1)=\frac{2m-1}{\sqrt{4m+1}}=\sqrt{4m+1}-\frac{m-1}{\frac{(m-1)\sqrt{4m+1}}{2m}}-\frac{2}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (22)
Hence the claim holds for $i=0,1$. It suffices to show that for any $i\in\{0\}\cup\llbracket m-2\rrbracket$, if $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Assume that $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then
$b:=\sqrt{4m+1}-\frac{m+1}{P(i)}=\sqrt{4m+1}-\frac{m+1}{\frac{2m-i}{\sqrt{4m+1}}}=\frac{(m-i-1)\sqrt{4m+1}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ (23)
By Claim 1, $\frac{(i+2)\sqrt{4m+1}}{2i+3}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$. Note that
$\sqrt{4m+1}-\frac{m-i-1}{b}-\frac{i+2}{\frac{(i+2)\sqrt{4m+1}}{2i+3}}$ $=$ $\sqrt{4m+1}-\frac{m-i-1}{\frac{(m-i-1)\sqrt{4m+1}}{2m-i}}-\frac{i+2}{\frac{(i+2)\sqrt{4m+1}}{2i+3}}$ (24)
$=$ $\frac{2m-i-2}{\sqrt{4m+1}},$
implying that $P(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$ by definition. Thus, Claim 3 holds.
By Lemma 4 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 4.
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for all integers $s$ in $[m,4m+2]$.
Proof. We first prove the following claims.
Claim 1: For any $i\in{\mathbb{N}}^{+}$, both $\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$ belong to ${\mathscr{W}}_{m+2}(\sqrt{4m+2})$.
Let $P(i)=\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $Q(i)=\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$. By Lemma 4, $P(1)=\frac{3m+1}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ and $Q(1)=\frac{3m+2}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 1 holds for $i=1$.
If $P(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$, then
$\sqrt{4m+2}-\frac{m+1}{\frac{(2i+1)m+i}{i\sqrt{4m+2}}}=\frac{\sqrt{4m+2}\,m(i+1)}{i(2m+1)+m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (25)
and
$Q(i+1)=\frac{(2(i+1)+1)m+(i+1)+1}{(i+1)\,\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m}{\frac{\sqrt{4m+2}\,m(i+1)}{i(2m+1)+m}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (26)
If $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$, then
$\sqrt{4m+2}-\frac{m}{\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}}=\frac{\sqrt{4m+2}\,(i+1)(m+1)}{i(2m+1)+m+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (27)
and
$P(i+1)=\frac{(2(i+1)+1)m+(i+1)}{(i+1)\,\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{\sqrt{4m+2}\,(i+1)(m+1)}{i(2m+1)+m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (28)
Thus, Claim 1 holds.
Claim 2: $\frac{2m-i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $i$ with $2\leq i\leq m+1$.
Let $W(i)=\frac{2m-i}{\sqrt{4m+2}}$. By Claim 1, $W(-2)=\frac{(2m+1)m+m}{m\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. By definition,
$W(-1)=\frac{2m+1}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{2m+2}{\sqrt{4m+2}}}=\sqrt{4m+2}-\frac{m+1}{W(-2)}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (29)
$W(0)=\frac{2m}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m+1}{\frac{2m+1}{\sqrt{4m+2}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (30)
and
$W(1)=\frac{2m-1}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{1}{\frac{2m+1}{\sqrt{4m+2}}}-\frac{m}{\frac{2m}{\sqrt{4m+2}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (31)
Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $2\leq i\leq 1$.
It suffices to show that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ implies that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Observe that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$,
$b:=\sqrt{4m+2}-\frac{m+1}{W(i)}=\sqrt{4m+2}-\frac{m+1}{\frac{2m-i}{\sqrt{4m+2}}}=\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}),$ (32)
and
$W(i+2)=\frac{2m-i-2}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m-i-1}{b}-\frac{i+2}{W(-1)},$ (33)
implying that $W(i+2)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 2 holds.
Claim 3: $\frac{s}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $s$ with $2m+1\leq s\leq 3m$.
By Claim 2, $\frac{2m+i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for $i=0,1$. For any $i\in\{0\}\cup\llbracket m-2\rrbracket$, it is shown in (32) that $\frac{(m-i-1)\sqrt{4m+2}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ Thus,
$\frac{2m+2+i}{\sqrt{4m+2}}=\sqrt{4m+2}-\frac{m-i-1}{\frac{(m-i-1)\cdot\sqrt{4m+2}}{2m-i}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2}).$ (34)
Hence Claim 3 holds.
By Lemma 4 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 5.
For any $m\in{\mathbb{N}}^{+}$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $s$ in $[m,4m+3]$.
Proof. We first prove the following claims.
Claim 1: For any $i\in{\mathbb{N}}^{+}$, $\frac{(i+1)\sqrt{4m+3}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ and $\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
Let $f_{1}(i)=\frac{(i+1)\sqrt{4m+3}}{2i+1}$ and $f_{2}(i)=\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}$. Obviously, $\sqrt{4m+3}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$,
$f_{2}(1)=\frac{3(m+1)}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ (35)
and
$f_{1}(1)=\frac{2\sqrt{4m+3}}{3}=\sqrt{4m+3}-\frac{m+1}{\frac{3(m+1)}{\sqrt{4m+3}}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (36)
Thus, Claim 1 holds for $i=1$.
Now assume that $t$ is an integer with $t\geq 2$ such that $f_{1}(i),f_{2}(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $i$ with $1\leq i\leq t-1$. Then, by definition,
$\frac{2mi+i+m}{i\sqrt{4m+3}}=\sqrt{4m+3}-\frac{i+1}{\frac{(i+1)\sqrt{4m+3}}{2i+1}}-\frac{m-i}{\frac{i\sqrt{4m+3}}{2i-1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (37)
$\sqrt{4m+3}-\frac{m+1}{\frac{2mi+i+m}{i\sqrt{4m+3}}}=\frac{(i+1)m\sqrt{4m+3}}{2mi+i+m}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (38)
$f_{2}(i+1)=\frac{(2i+3)(m+1)}{(i+1)\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m}{\frac{(i+1)m\sqrt{4m+3}}{2mi+i+m}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (39)
and
$f_{1}(i+1)=\frac{(i+2)\sqrt{4m+3}}{2i+3}=\sqrt{4m+3}-\frac{m+1}{f_{2}(i+1)}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (40)
Thus, Claim 1 holds.
Claim 2: For any integer $i$ with $3\leq i\leq m$, $\frac{2m-i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
Let $g(i)=\frac{2m-i}{\sqrt{4m+3}}$. By Claim 1, $g(-3)=\frac{2m+3}{\sqrt{4m+3}}=f_{2}(m+1)$. By definition,
$\sqrt{4m+3}-\frac{m+1}{f_{1}(m)}=\sqrt{4m+3}-\frac{m+1}{\frac{(m+1)\sqrt{4m+3}}{2m+1}}=\frac{2(m+1)}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (41)
$\sqrt{4m+3}-\frac{m+1}{\frac{2(m+1)}{\sqrt{4m+3}}}=\frac{\sqrt{4m+3}}{2}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (42)
$\sqrt{4m+3}-\frac{m+1}{\frac{\sqrt{4m+3}}{2}}=\frac{2m+1}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (43)
$\sqrt{4m+3}-\frac{m+1}{\frac{2m+1}{\sqrt{4m+3}}}=\frac{m\sqrt{4m+3}}{2m+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (44)
and
$\sqrt{4m+3}-\frac{m}{\frac{m\sqrt{4m+3}}{2m+1}}-\frac{1}{\frac{\sqrt{4m+3}}{2}}=\frac{2m}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (45)
implying that $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each $i\in\{-2,-1,0\}$.
It remains to show that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$, $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ implies that $g(i+1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Observe that
$\sqrt{4m+3}-\frac{m+1}{g(i)}=\sqrt{4m+3}-\frac{(m+1)}{\frac{2m-i}{\sqrt{4m+3}}}=\frac{(m-i-1)\sqrt{4m+3}}{2m-i}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (46)
By (42), $\frac{\sqrt{4m+3}}{2}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Thus,
$g(i+1)=\frac{2m-i-1}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{m-i-1}{\frac{(m-i-1)\sqrt{4m+3}}{2m-i}}-\frac{i+2}{\frac{\sqrt{4m+3}}{2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}),$ (47)
Hence Claim 2 follows.
Claim 3: For any $s\in{\mathbb{N}}^{+}$ with $2m\leq s\leq 4m+3$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
By Lemma 4, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $3m+2\leq s\leq 4m+3$.
Let $h(i)=\frac{2m+i}{\sqrt{4m+3}}$. By Claim 2, $h(0),h(1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. For any $i\in\{0\}\cup\llbracket m\rrbracket$, by Claim 1, $\frac{(m-i+1)\sqrt{4m+3}}{2(m-i)+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$, implying that
$h(i+2)=\frac{2m+2+i}{\sqrt{4m+3}}=\sqrt{4m+3}-\frac{(m+1-i)}{\frac{(m-i+1)\sqrt{4m+3}}{2\cdot(m-i)+1}}-\frac{i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3}).$ (48)
Thus, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $2m\leq s\leq 3m+1$. Hence Claim 3 holds. The result then follows from Claims 2 and 3. $\Box$
We are now going to prove Theorem 3.
Proof of Theorem 3: The necessity follows from Theorem 1. Now we prove the sufficiency.
Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. By Lemma 4 and Propositions 234 and 5, for each integer $s$ with $r_{0}\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})\subseteq{\mathscr{W}}_{s}(\sqrt{k})$ holds. Note that
$s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}.$ (49)
By Corollary 3, for any integer $s$ in $r_{0},k]$, there exists a tree $T$ with $\Delta(T)=s$ and $\rho(T)=\sqrt{k}$. $\Box$
5 Further study
Recall that $\mathbb{P}$ is the set of $\rho(T)$ over all trees $T$ of order at least $2$. Problem 1 asks an interesting question on the existence of a tree $T$ with $\rho(T)=\alpha$ and $\Delta(T)=r$ for any given $\alpha\in\mathbb{P}\setminus\{1\}$ and integer $r$ satisfying $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
Another interesting problem is on the numbers contained in $\mathbb{P}$. Let $\cal P$} be the set of real-rooted monic polynomials with integer coefficients. Salez [28] showed that every totally real algebraic integer (i.e. a root of some polynomial in $\cal P$}) is an eigenvalue of some finite tree. Now let $\mathbb{P}^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
Problem 2.
Is there a positive real number $c$ such that $\{\alpha\in\mathbb{P}^{\prime}:\alpha\geq c\}\subseteq\mathbb{P}$?
It can be proved by applying Theorem 2 that, for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ if 4|(pq)4|(p-q), $\sqrt{p+\sqrt{q}}\in\mathbb{P}$ if $p\geq q$, and $\frac{\sqrt{(2p+1)+\sqrt{4q+1}}}{\sqrt{2}}\in\mathbb{P}$ if $q\leq p^{2}+p$. As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ whenever 4|(pq)4|(p-q).
Proposition 6.
For any $p,q\in{\mathbb{N}}^{+}$ with $p\geq q$, if pqp-q is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\left\{q,\frac{p-q}{4}\right\}$ and $\rho(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
Proof. Assume that pq=4dp-q=4d. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
$\sqrt{q}=\frac{\sqrt{p}+\sqrt{q}}{2}-\frac{d}{\frac{\sqrt{p}+\sqrt{q}}{2}}\in{\mathscr{W}}_{r}(\alpha)$ (50)
and
$\frac{\sqrt{p}-\sqrt{q}}{2}=\frac{\sqrt{p}+\sqrt{q}}{2}-\frac{q}{\sqrt{q}}\in{\mathscr{W}}_{r}(\alpha).$ (51)
Observe that
$d\cdot\left(\frac{\sqrt{p}-\sqrt{q}}{2}\right)^{-1}=\frac{\frac{p-q}{4}}{\frac{\sqrt{p}-\sqrt{q}}{2}}=\frac{\sqrt{p}+\sqrt{q}}{2}=\alpha.$ (52)
By Corollary 4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\rho(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
We wonder if 4|(pq)4|(p-q) is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\mathbb{P}$.
Problem 3.
Is it true that for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ if and only if $4\big{|}(p-q)$?
Let $T$ be any tree with $\alpha=\rho(T)$ and $r=\Delta(T)$. Then, Theorem 1 yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$. From the proving of Lemma 1, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
Problem 4.
Let $\lambda\in\mathbb{P}$ and $r\in{\mathbb{N}}^{+}$ with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
Note that some similar work on Laplacian spectral radii of trees has been completed in [8].
Conflict of interest: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability statements: Not applicable.
Acknowledgement
This research is supported by NSFC (No. 12101347 and 12371340), NSF of Shandong Province (No. ZR2021QA085) and the Ministry of Education, Singapore, under its Academic Research Tier 1 (RG19/22). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.